OpenStudy (geekfromthefutur):

The conical hat shown here is a traditional sun-protective straw hat originating in eastern Asia. What is the approximate outer surface area of one of these hats that has a radius of 9 inches and a slant height of 12 inches? Keep in mind that there is an opening in the hat where the base surface would normally be. Show the steps of your solution and explain your work.

OpenStudy (ashleyisakitty):

Try to do this: pi * radius * slant

OpenStudy (ashleyisakitty):

pi * 9 * 12 = ?

OpenStudy (geekfromthefutur):

108

OpenStudy (geekfromthefutur):

im not a big fan of math sometimes.

OpenStudy (ashleyisakitty):

correct!

OpenStudy (geekfromthefutur):

:D

OpenStudy (ashleyisakitty):

Wait! My steps were a bit off I think. It may be pi*9*12=96*pi=301.